# Conformalized Polynomial Chaos Expansion for Uncertainty-aware Surrogate Modeling

### v1

@@ 4.  Summary, conclusions, and outlook
This work presented a new method that combines regression-based PCEs with jackknife-based CP for constructing uncertainty-aware, data-driven surrogate models. Using this method, the data-driven PCE is complemented with a predictive interval that quantifies the uncertainty in its predictions. Crucially, this predictive interval is estimated with very low computational overhead, by leveraging the linearity of the PCE regression model to compute LOO residuals and LOO predictions via analytical, closed-form expressions. The latter are used by the jackknife and jackknife+ CP methods to estimate the predictive interval.
@@
The conformalized PCE method was applied to four different benchmark models, with one, six, seven, and ten input dimensions, respectively. Total-degree PCEs were employed for the multivariate models. Different PCE configurations were used, by varying the maximum polynomial degree, experimental design size, CP method (jackknife versus jackknife+), and non-conformity score type (normalized versus non-normalized). Based on the numerical results, the conformalized PCE is able to provide empirical coverages close to the target level. Training data availability, equivalently, experimental design size is important for reducing the variability in predictive interval width and coverage and stabilizing the PCE. The same can be observed when increasing the PCE maximum polynomial degree, thus increasing its accuracy as well (assuming again sufficient training data). The jackknife method typically results in more conservative conformalized PCEs, with wider predictive intervals and empirical coverages above the target level. Contrarily, the jackknife+ method yields coverages closer to the target level. Non-conformity score normalization does not seem to have a significant impact on predictive interval width and coverage, at least for the examined benchmark models.
@@
The conformalized PCE method developed in this work shows definite promise, however, it can be improved in several aspects. First, this work has considered total-degree PCEs only, thus limiting the selected test-cases to low dimensions and/or polynomial degrees. The integration of CP within sparse and adaptive PCEs would allow to extend the method to more challenging problem settings. Second, in its current form, the conformalized PCE method is applicable to scalar outputs only. In principle, it can be applied elementwise to address multidimensional outputs. However, this would be a suboptimal approach, as it does not account for correlations among the output’s elements and would incur high computational costs with increasing output dimension. The development of a conformalized PCE method suitable for multivariate outputs seems to be a natural follow-up, for example based on the combination of multi-output PCEs [51, 52, 53] and multi-output CP [54, 55]. We aim to address these limitations in dedicated follow-up works.
@@
Appendix A Leave-one-out residuals in closed-form
@@
Appendix B Leave-one-out predictions in closed-form

### v2

@@ 4.  Summary, conclusions, and outlook
This work presented a new method that combines regression-based PCEs with jackknife-based CP for constructing uncertainty-aware, data-driven surrogate models. Using this method, data-driven PCEs are complemented with predictive intervals that quantify the uncertainty in their predictions. The predictive intervals are estimated with very low computational overhead, by leveraging the linearity of the PCE regression model. In particular, analytical, closed-form expressions can be used for computing the LOO residuals and LOO predictions employed by the jackknife and jackknife+ CP methods for predictive interval estimation.
@@
The conformalized PCE method was demonstrated on four benchmark models and two electrical engineering design applications concerning a power electronics heat sink and a high-gradient-field electromagnet. Different conformalized PCE configurations were examined, based on maximum polynomial degree, experimental design size, CP method (jackknife/jackknife+), and non-conformity score type (normalized/non-normalized). Based on the numerical results, the conformalized PCE is in general able to provide empirical coverages close to the required target level. Training data availability, equivalently, experimental design size is a most important factor for providing the required coverage with the minimum possible interval width. Moreover, considering PCE models trained with different datasets, larger training datasets result in reduced variations in coverage and predictive interval width. The jackknife method typically results in more conservative predictive uncertainty estimates, with wider intervals and empirical coverages above the target level. The jackknife+ method yields coverages closer to the target level. Non-conformity score normalization does not seem to have a significant or consistent impact on predictive interval width and coverage, at least for the examined benchmark models and engineering use-cases.
@@
The conformalized PCE method developed in this work shows definite promise as uncertainty-aware, data-driven, surrogate modeling method. Nonetheless, it can be improved in several aspects. First, this work has considered total-degree PCEs only, thus limiting the selected benchmark models and engineering use-cases to low dimensions and/or polynomial degrees. The integration of CP within sparse PCE algorithms [luethen2021sparse, luethen2022automatic] would allow to extend the method to more challenging problem settings. Second, in its current form, the conformalized PCE method is applicable to scalar outputs only. In principle, it can be applied elementwise to address multidimensional outputs. However, this would be a suboptimal approach, as it does not account for correlations among the output’s elements and would incur high computational costs with increasing output dimension. The development of a conformalized PCE method suitable for multivariate outputs seems to be a natural follow-up, for example based on the combination of multi-output PCEs [hawchar2017principal, bhattacharyya2020uncertainty, loukrezis2025multivariate] and multi-output CP [zhang2023improved, xu2023conformal]. We aim to address these limitations in dedicated follow-up works.
Acknowledgements
Dimitris G. Giovanis was supported by the Department of Energy (DOE) under Grant No. DE-SC0024162.
@@ Acknowledgements
Appendix A Leave-one-out residuals in closed form
@@
Appendix B Leave-one-out predictions in closed form